30g of NaOH was dissolved in enough water to produce a 125ml solution.The temperature increased from 22°C to 79°C.The density of the solution is 10g/ml.The specific heat capacity of the solution is 4.18J/gC.What is the enthalpy change for the dissolution of NaOH?

1
Expert's answer
2022-01-10T13:57:55-0500

∆T = 79-22= 57

q=4.18×57×30= 7147.8J